How do I wrtie an equation that will be perpendicular to 3x-5y=10 and passes through (-6,2)?
Answer by stanbon(57397)

How do I wrtie an equation that will be perpendicular to 3x-5y=10 and passes through (-6,2)?
-------------
slope of the given line is 3/5
----------------
Therefore slope of the perpendicular line is -5/3
----------------
Find the y-intercept if y = mx+b:
2 = (-5/3)(-6) + b
2 = 10 + b
b = -8
---------------------------
Equation of the line:
y = (-5/3)x - 8
============================
Cheers,
stan H.